# A Starting Point for Dynamic Community Detection with Leiden Algorithm

### v1

@@ 4.1.1. System
In our experiments, we employ a server equipped with an x86-based 64-bit AMD EPYC-7742 processor. It operates at $2.25$ GHz and is paired with $512$ GB of DDR4 system memory. Each core features a $4$ MB L1 cache, a $32$ MB L2 cache, and a shared L3 cache of $256$ MB. The server runs on Ubuntu 20.04.
@@ 4.1.2. Configuration
We utilize a 32-bit unsigned integer format for vertex IDs and 32-bit floating-point representation for edge weights. However, for hashtable values, total edge weight calculations, modularity computation, and any instances requiring aggregation or summation of floating-point values, we employ the 64-bit floating-point format. Affected vertices are denoted by an 8-bit integer vector. In computing the weighted degree of each vertex, the local moving phase, the refinement phase, and aggregating edges for the super-vertex graph, we leverage OpenMP’s dynamic schedule with a chunk size of $2048$ to facilitate dynamic workload balancing among threads. Key parameters are configured as follows: the iteration tolerance $\tau$ is set to $10^{-2}$, the tolerance drop per pass $TOLERANCE\_DECLINE\_FACTOR$ to $10$ (as part of threshold-scaling optimization), the maximum number of iterations per pass $MAX\_ITERATIONS$ to $20$, and the maximum number of passes $MAX\_PASSES$ to $10$. Additionally, the aggregation tolerance $\tau_{agg}$ is set to $0.8$ for large (static) graphs with randomly generated batch updates, while it remains disabled (i.e., $\tau_{agg}$ set to $1$) for real-world dynamic graphs. Unless explicitly specified, all parallel implementations are executed with a default of $64$ threads, matching the available number of cores on the system. Compilation is carried out using GCC 9.4 and OpenMP 5.0 (OpenMP Architecture Review Board, 2018).
4.1.3. Dataset
To conduct experiments on large static graphs with random batch updates, we employ $12$ graphs listed in Table 1, obtained from the SuiteSparse Matrix Collection (Kolodziej et al., 2019). For these graphs, the number of vertices range from $3.07$ to $214$ million, and the number of edges range from $25.4$ million to $3.80$ billion. For experiments involving real-world dynamic graphs, we utilize five temporal networks sourced from the Stanford Large Network Dataset Collection (Leskovec and Krevl, 2014), detailed in Table 2. Here, the number of vertices span from $24.8$ thousand to $2.60$ million, temporal edges from $507$ thousand to $63.4$ million, and static edges from $240$ thousand to $36.2$ million. With each graph, we ensure that all edges are undirected and weighted, with a default weight of $1$.
Table 1. List of $12$ graphs retrieved from the SuiteSparse Matrix Collection (Kolodziej et al., 2019) (with directed graphs indicated by *). Here, $|V|$ denotes the number of vertices, $|E|$ denotes the number of edges (after making the graph undirected by adding reverse edges), and $|\Gamma|$ denotes the number of communities obtained with Static Leiden algorithm (Sahu, 2023).
@@ 4.1.4. Batch generation
For experiments on large graphs with random batch updates, we take each base graph from Table 1 and generate random batch updates (Zarayeneh and Kalyanaraman, 2021), comprising an $80\%:20\%$ mix of edge insertions and deletions to emulate realistic batch updates, with each edge having a weight of $1$. To prepare the set of edges for insertion, we select vertex pairs with equal probability. For edge deletions, we uniformly delete existing edges. To simplify, no new vertices are added or removed from the graph. The batch size is measured as a fraction of the edges in the original graph, ranging from $10^{-7}$ to $0.1$ (i.e., $10^{-7}|E|$ to $0.1|E|$). Multiple batches are generated for each size to allow averaging. All batch updates are undirected, meaning for every edge insertion $(i,j,w)$ in the batch update, the edge $(j,i,w)$ is also included. We employ five distinct random batch updates for each batch size and report the average across these runs in our experiments. For experiments involving real-world dynamic graphs, we start by loading $90\%$ of each graph from Table 2, and ensure all edges are weighted with a default weight of $1$ and are undirected by adding reverse edges. Subsequently, we load $B$ edges in $100$ batch updates. Here, $B$ denotes the desired batch size, specified as a fraction of the total number of temporal edges $|E_{T}|$ in the graph, and ensure the batch update is undirected.
@@ 4.1.5. Measurement
We evaluate the runtime of each approach on the entire updated graph, including the local-moving phase, refinement phase, aggregation phase, initial and incremental marking of affected vertices, convergence detection, and all necessary intermediary steps, but exclude memory allocation and deallocation time. We assume that the total edge weight of the graphs is known and can be tracked with each batch update.
@@
Figure 3. Runtime (logarithmic scale) of our multicore implementation of Static, Naive-dynamic (ND), Delta-screening (DS), and Dynamic Frontier (DF) Leiden on large (static) graphs with randomly generated batch updates. The size of these batch updates ranges from $10^{-7}|E|$ to $0.1|E|$ in multiples of $10$, with the updates comprising $80\%$ edge insertions and $20\%$ edge deletions to simulate realistic dynamic graph changes. The right subfigure shows the runtime of each algorithm for individual graphs in the dataset, while the left subfigure displays overall runtimes using the geometric mean for consistent scaling across graphs. Furthermore, the speedup of each algorithm compared to Static Leiden is indicated on the respective lines.
@@
Figure 4. Modularity comparison of our multicore implementation of Static, Naive-dynamic (ND), Delta-screening (DS), and Dynamic Frontier (DF) Leiden on large (static) graphs with randomly generated batch updates. The size of these batch updates ranges from $10^{-7}|E|$ to $0.1|E|$ in multiples of $10$, with the updates comprising $80\%$ edge insertions and $20\%$ edge deletions to simulate realistic dynamic graph changes. The right subfigure shows the modularity for each algorithm for individual graphs in the dataset, while the left subfigure displays overall modularity using arithmetic mean.
(a) Overall Runtime
(b) Overall Modularity of communities obtained
(c) Runtime on each dynamic graph
(d) Modularity in communities obtained on each dynamic graph
@@ 4.2. Performance Comparison
4.2.1. Results on large graphs with random batch updates
We now evaluate the performance of our parallel implementations of Static, Naive-dynamic (ND), Delta-screening (DS), and Dynamic Frontier (DF) Leiden on large (static) graphs listed in Table 1, using randomly generated batch updates. As detailed in Section 4.1.4, the batch updates range in size from $10^{-7}|E|$ to $0.1|E|$ (in multiples of $10$), consisting of $80\%$ edge insertions and $20\%$ edge deletions to simulate realistic scenarios. Reverse edges are included with each batch update to maintain the graph as undirected. As stated in Section 4.1.4, we generate $5$ different random batch updates for each batch size to reduce measurement noise. Figure 3 presents the runtime of Static, ND, DS, and DF Leiden, while Figure 4 shows the modularity of the communities obtained with each approach.
Figure 3(a) demonstrates that ND, DS, and DF Leiden achieve mean speedups of 1.25×1.25$\times$1.25 ×, 1.24×1.24$\times$1.24 ×, and 1.37×1.37$\times$1.37 ×, respectively, when compared to Static Leiden. Let us now discuss why ND, DS, and DF Leiden exhibit only modest speedups over Static Leiden. Unlike Louvain, we must always run the refinement phase of Leiden algorithm to avoid internally disconnected communities. Nonetheless, the refinement phase splits the communities obtained/updated from the local-moving phase into several smaller sub-communities. Stopping the passes early, as done with DF Louvain, leads to low modularity for ND/DS/DF Leiden because sufficiently large high-quality clusters have not yet formed due to the refinement phase. Further, ND/DS/DF Leiden can only reduce the runtime of the local-moving phase of the first pass of the Leiden algorithm. However, only about $37\%$ of the runtime in Static Leiden is spent in the local-moving phase of the first pass. These factors constrain the speedup potential of ND, DS, and DF Leiden over Static Leiden. Figures 4(a) and 4(b) show that ND, DS, and DF Leiden achieve communities with approximately the same modularity as Static Leiden. Thus, for large graphs with random batch updates, DF Leiden appears to be the dynamic community detection method.
@@ 4.2.1. Results on large graphs with random batch updates
Also note in Figure 3 that the runtime of Static Leiden increases with larger batch updates. This phenomenon can be attributed to the random batch updates, which indiscriminately disturb the original community structure, necessitating Static Leiden to undergo more iterations to converge, rather than solely due to the increased number of edges in the graph.
4.2.2. Results on real-world dynamic graphs
We also evaluate the performance of our parallel implementations of Static, ND, DS, and DF Leiden on real-world dynamic graphs listed in Table 2. These evaluations are performed on batch updates ranging from $10^{-5}|E_{T}|$ to $10^{-3}|E_{T}|$ in multiples of $10$. For each batch size, as described in Section 4.1.4, we load $90\%$ of the graph, add reverse edges to ensure all edges are undirected, and then load $B$ edges (where $B$ is the batch size) consecutively in $100$ batch updates. Figure 5(a) shows the overall runtime of each approach across all graphs for each batch size, while Figure 5(b) depicts the overall modularity of the obtained communities. Additionally, Figures 5(c) and 5(d) present the mean runtime and modularity of the communities obtained with each approach on individual dynamic graphs in the dataset.
Figure 5(a) illustrates that ND Leiden is, on average, 1.14×1.14$\times$1.14 × faster than Static Leiden for batch updates ranging from $10^{-5}|E_{T}|$ to $10^{-3}|E_{T}|$. In comparison, DS and DF Leiden exhibit average speedups of $1\times1$ and 1.09×1.09$\times$1.09 ×, respectively, over Static Leiden for the same batch updates. We now explain why ND, DS, and DF Leiden achieve only minor speedups over Static Leiden. Unlike with large graphs and random batch updates, on real-world dynamic graphs, we disable the aggregation tolerance $\tau_{agg}$. Our experiments indicate that disabling aggregation tolerance allows for higher modularity but increases the number of passes performed. This results in a reduced portion of runtime spent in the local-moving phase of the first pass, which is the phase that can be minimized by the dynamic approaches. Specifically, we observe that only about $20\%$ of the overall runtime of Static Leiden is spent in the local-moving phase of the first pass. Consequently, the speedup achieved by ND, DS, and DF Leiden relative to Static Leiden is limited. Furthermore, our observations suggest that while DF Leiden can reduce the time spent in the local-moving phase slightly more than ND Leiden, it incurs increased runtimes in the refinement and aggregation phases of the algorithm. This is likely because ND Leiden is able to optimize clusters more effectively than DF Leiden, which accelerates the refinement and aggregation phases for ND Leiden, albeit by a small margin. This explains the lower speedup of DF Leiden compared to ND Leiden on real-world dynamic graphs.
@@ 5. Conclusion
In conclusion, this report extended three approaches for community detection on dynamic graphs — Naive-dynamic (ND), Delta-screening (DS), and Dynamic Frontier (DF) — to our multicore implementation (Sahu, 2023) of the Leiden algorithm (Traag et al., 2019), an algorithm renowned for its robust community detection capabilities. Our experiments conducted on a server featuring a 64-core AMD EPYC-7742 processor reveal that ND, DS, and DF Leiden achieve speedups of 1.25×1.25$\times$1.25 ×, 1.24×1.24$\times$1.24 ×, and 1.37×1.37$\times$1.37 × respectively, when applied to large graphs with randomized batch updates, compared to Static Leiden. However, when applied to real-world dynamic graphs, ND, DS, and DF Leiden demonstrate speedups of 1.14×1.14$\times$1.14 ×, $1\times1$, and 1.09×1.09$\times$1.09 × respectively. The relatively modest speedups are due to the dynamic approaches being capable of reducing the runtime of only the local-moving phase of the initial pass of the Leiden algorithm, while still necessitating the execution of all subsequent passes akin to Static Leiden, in order to attain high-modularity communities devoid of internal disconnections. We hope these preliminary findings motivate further exploration of dynamic approaches to the Leiden algorithm.

### v4

@@ 5.1.1. System
We employ a server equipped with an x86-based 64-core AMD EPYC-7742 processor. It operates at a clock frequency of $2.25$ GHz and is paired with $512$ GB of DDR4 system memory. Each core features a $4$ MB L1 cache, a $32$ MB L2 cache, and a shared L3 cache of $256$ MB. The server runs on Ubuntu 20.04.
@@ 5.1.2. Configuration
Table 2. List of $12$ graphs retrieved from the SuiteSparse Matrix Collection (Kolodziej et al., 2019) (with directed graphs indicated by *). Here, $|V|$ denotes the number of vertices, $|E|$ denotes the number of edges (after making the graph undirected by adding reverse edges), and $|\Gamma|$ denotes the number of communities obtained with Static Leiden algorithm (Sahu et al., 2024a).
@@
We utilize 32-bit unsigned integers for vertex IDs and 32-bit floating-point for edge weights. For computation involving floating-point aggregation, like total edge weights and modularity, we switch to 64-bit floating-point. Affected vertices and changed communities are represented by 8-bit integer vectors. During the local-moving, refinement, and aggregation phases, we employ OpenMP’s dynamic scheduling with a chunk size of $2048$ for load balancing. However, for the aggregation phase in Naive-dynamic (ND), Delta-screening (DS), and Dynamic Frontier (DF) Leiden, a chunk size of $32$ is used. We utilize state-of-the-art parallel Leiden (Sahu et al., 2024a), incorporating our ND, DS, and DF Leiden atop it. For real-world dynamic graphs (discussed in Section A.5), which aggregate slowly, we disable aggregation tolerance - setting it to $1$. This ensures all desirable community aggregations occur, yielding higher modularity scores compared to the default of $0.8$. For consistency, we apply this to static Leiden as well - to avoid comparing apples-to-oranges. We run all implementations on $64$ threads and compile using GCC 9.4 with OpenMP 5.0.
5.1.3. Dataset
We conduct experiments on $12$ large real-world graphs with random batch updates, listed in Table 2, obtained from the SuiteSparse Matrix Collection. These graphs have between $3.07$ million and $214$ million vertices, and $25.4$ million to $3.80$ billion edges. For experiments involving real-world dynamic graphs, we utilize five temporal networks sourced from the Stanford Large Network Dataset Collection (Leskovec and Krevl, 2014), detailed in Table 3. Here, the number of vertices span from $24.8$ thousand to $2.60$ million, temporal edges from $507$ thousand to $63.4$ million, and static edges from $240$ thousand to $36.2$ million. However, we would like to note that existing temporal graphs in the SNAP repository (Leskovec and Krevl, 2014) are mostly small, limiting their utility for studying our proposed parallel algorithms. With each graph, we ensure that all edges are undirected and weighted, with a default weight of $1$. Since most publicly available real-world weighted graphs are small in size, we do not use them here — although our parallel algorithms can handle weighted graphs without any modification.
We also do not use SNAP datasets with ground-truth communities, as they are non-disjoint, while our work focuses on disjoint communities. Nevertheless, we wish to note that the goal of community detection is not necessarily to match the ground truth, as this may reflect a different aspect of network structure or be unrelated to the actual structure, and focusing solely on this correlation risks overlooking other meaningful community structures (Peel et al., 2017).
@@ 5.1.4. Batch generation
We take each base graph from Table 2 and generate random batch updates (Zarayeneh and Kalyanaraman, 2021), comprising an $80\%:20\%$ mix of edge insertions and deletions to emulate realistic batch updates, with each edge having a weight of $1$. Batch updates are generated with unit-weighted edges for simplicity, without affecting results, as the algorithms handle arbitrary edge weights. Since the random batch updates are processed separately, the graph does not converge toward a random structure. This enables testing on diverse graph classes. All updates are undirected; for every edge insertion $(i,j,w)$, we also include $(j,i,w)$. To prepare the set of edges for insertion, we select vertex pairs with equal probability. For edge deletions, we uniformly delete existing edges. To simplify, no new vertices are added or removed from the graph. The batch size is measured as a fraction of the edges in the original graph, ranging from $10^{-7}$ to $0.1$ (i.e., $10^{-7}|E|$ to $0.1|E|$). For a billion-edge graph, this translates to a batch size ranging from $100$ to $100$ million edges. We employ five distinct random batch updates for each batch size and report the average across these runs. Note that dynamic graph algorithms are helpful for small batch updates in interactive applications. For large batches, it is usually more efficient to run the static algorithm.
@@ 5.1.5. Measurement
We evaluate the runtime of each method on the entire updated graph, encompassing all stages: the local-moving phase, refinement phase, aggregation phase, initial and incremental marking of affected vertices, convergence detection, and any necessary intermediate steps. To mitigate the impact of noise, we follow the standard practice of running each experiment multiple times. We assume that the total edge weight of the graphs is known and can be tracked with each batch update. As our baseline, we use the most efficient multicore implementation of Static Leiden (Sahu et al., 2024a), which has been shown to outperform even GPU-based implementations. Given that modularity maximization is NP-hard, all existing polynomial-time algorithms are heuristic; thus, we assess the optimality of our dynamic algorithms by comparing their convergence to the modularity score achieved by the static algorithm. Finally, none of the analyzed algorithms — Static, ND, DS, and DF Leiden — produce communities that are internally disconnected, so we omit this from our figures.
@@
Figure 7. Runtime (logarithmic scale) of our multicore implementation of Naive-dynamic (ND), Delta-screening (DS), and Dynamic Frontier (DF) Leiden, compared to Static Leiden (Sahu et al., 2024a), on large graphs with randomly generated batch updates. The size of these batch updates ranges from $10^{-7}|E|$ to $0.1|E|$ in multiples of $10$, with the updates comprising $80\%$ edge insertions and $20\%$ edge deletions to simulate realistic dynamic graph changes. The right subfigure shows the runtime of each algorithm for individual graphs in the dataset, while the left subfigure displays overall runtimes using the geometric mean for consistent scaling across graphs. The speedup of each algorithm compared to Static Leiden is indicated on the respective lines.
@@
Figure 8. Modularity comparison of our multicore implementation of Naive-dynamic (ND), Delta-screening (DS), and Dynamic Frontier (DF) Leiden, compared to Static Leiden (Sahu et al., 2024a), on large graphs with randomly generated batch updates. These batch updates vary in size from $10^{-7}|E|$ to $0.1|E|$ in powers of 10, consisting of $80\%$ edge insertions and $20\%$ edge deletions to mimic realistic dynamic graph updates. Here, the right subfigure shows the modularity for each algorithm for individual graphs in the dataset, while the left subfigure displays overall modularity obtained using arithmetic mean.
@@ 5.2. Performance Comparison
We now evaluate the performance of our parallel implementations of ND, DS, and DF Leiden against Static Leiden (Sahu et al., 2024a), on large graphs with randomly generated batch updates. Static Leiden is rerun from scratch after each update, whereas ND, DS, and DF Leiden begin with prior community assignments. Additionally, DS and DF Leiden identify affected vertices, processing only these during local-moving phase. Our methods are applicable to ND, DS, and DF Leiden. Original Leiden (Traag et al., 2019) is sequential; thus we compare against the fastest parallel Leiden (Sahu et al., 2024a). As detailed in Section 5.1.4, these updates range in size from $10^{-7}|E|$ to $0.1|E|$, with $80\%$ edge insertions and $20\%$ edge deletions. Each batch update includes reverse edges to keep the graph undirected. Figure 7(b) presents the execution time of each algorithm across individual graphs, while Figure 7(a) depicts overall runtimes using geometric mean for consistent scaling across differing graph sizes. In addition, Figure 8(b) shows the modularity of the algorithms for individual graphs in the dataset, and Figure 7(a) presents the overall modularity of each of the algorithms, calculated using arithmetic mean.
From Figure 7(a), we observe that ND, DS, and DF Leiden achieve mean speedups of 1.37×1.37$\times$1.37 ×, 1.47×1.47$\times$1.47 ×, and 1.98×1.98$\times$1.98 ×, respectively, when compared to Static Leiden. This speedup is higher on smaller batch updates, with ND, DS, and DF Leiden being on average 1.46×1.46$\times$1.46 ×, 1.74×1.74$\times$1.74 ×, and 3.72×3.72$\times$3.72 ×, respectively, when compared to Static Leiden, on batch updates of size $10^{-7}|E|$. Figure 7(b) shows that ND, DS, and DF Leiden offer moderate speedups over Static Leiden on web graphs and social networks — while on road networks (with low average degree), ND Leiden marginally outperforms Static Leiden, and DS and DF Leiden perform significantly better. This is due to DS and DF Leiden processing significantly fewer affected vertices/communities on road networks. On protein k-mer graphs, DS Leiden offers only a slight improvement. However, on small batch updates, DF Leiden outperforms Static Leiden by a large margin.
In terms of modularity, as Figures 8(a) and 8(b) show, ND, DS, and DF Leiden achieve communities with approximately the same modularity as Static Leiden. However, on social networks, ND, DS, and DF Leiden moderately outperform Static Leiden. This is likely due to social networks not having a strong enough community structure, and thus requiring more iterations to converge to a better community assignment. On such graphs, Static Leiden fails to attain better modularity due to it limiting the number of iterations per pass performed, in favor of faster runtimes. However, since ND, DS, and DF Leiden do not start from scratch, but leverage prior community assignments, they are able to converge faster, and with higher modularity than Static Leiden (Figure 8) — particularly on social networks, where poor community structures make clustering more challenging. Thus, on average, the modularity of communities from Static Leiden is slightly lower. However, we notice that modularity of communities identified by our algorithms do not differ by more than $0.002$ from Static Leiden, on average.
@@
Also note in Figure 7 that the runtime of Static Leiden increases with larger batch updates. This effect is primarily due to two factors: (1) Larger batches increase graph size, as $80\%$ of updates are edge insertions; and (2) Random edge updates disrupt original community structure, requiring more iterations to converge. This also accounts for the observed decline in modularity of the algorithms with larger batch sizes, as seen in Figure 8.
Let us now discuss why ND, DS, and DF Leiden exhibit only modest speedups over Static Leiden. Unlike Louvain algorithm, we must always run the refinement phase of Leiden algorithm to avoid badly connected and internally disconnected communities. Nonetheless, the refinement phase splits the communities obtained/updated from the local-moving phase into several smaller sub-communities. Stopping the passes early leads to low modularity scores for ND/DS/DF Leiden because sufficiently large high-quality clusters have not yet formed. Further, ND/DS/DF Leiden can only reduce the runtime of the local-moving phase of the first pass of the Leiden algorithm. However, only about $37\%$ of the runtime in Static Leiden is spent in the local-moving phase of the first pass. These factors constrain the speedup potential of ND, DS, and DF Leiden over Static Leiden. Regardless, on large graphs with random batch updates, DF Leiden appears to be the dynamic community detection method of choice for tracking evolving communities. Performance comparison on real-world dynamic graphs is given in Section A.5.
5.3. Affected vertices and Changed communities
We now study the fraction of vertices marked as affected, and the communities marked as changed (to keeping track of the subset of communities to be refined), by DS and DF Leiden on instances from Table 2, with batch updates ranging from $10^{-}7|E|$ to $0.1|E|$. This tracking occurs only in the first pass of the Leiden algorithm, as detailed in Sections 4.3 and 4.4, and is shown in Figure 9.
Figure 9. Fraction of vertices marked as affected (dotted lines), and communities marked as changed (solid lines) with Delta-screening (DS) and Dynamic Frontier (DF) Leiden on graphs in Table 2. The labels indicate the ratio of communities marked as changed by DS Leiden to that of DF Leiden.
Figure 10. Strong scalability of ND, DS, and DF Leiden, compared to Static Leiden, on batch updates of size $10^{-3}|E|$. The number of threads is doubled from $1$ to $64$ (logarithmic scale).
We notice from Figure 9 that DF Leiden marks fewer affected vertices and changed communities than DS Leiden, but their runtime difference is smaller, as many vertices marked by DS Leiden do not change their community labels, and converge quickly. However, as expected, performance of the algorithms decline with more affected vertices, as Figures 7 and 9 show. Further, while a high number of changed communities increases refinement costs in the first pass, the cost of subsequent passes only depends on the graph’s nature.
5.4. Scalability
Finally, we study the strong-scaling behavior of ND, DS, and DF Leiden, and compare it with Static Leiden. For this, we fix the batch size at $10^{-3}|E|$, vary the threads from $1$ to $64$, and measure the speedup of each algorithm with respect to its sequential version.
As shown in Figure 10, ND, DS, and DF Leiden obtain a speedup of 10.2×10.2$\times$10.2 ×, $\times9$, and 9.0×9.0$\times$9.0 × respectively at $32$ threads (with respect to sequential); with their speedup increasing at a mean rate of 1.59×1.59$\times$1.59 ×, 1.58×1.58$\times$1.58 ×, and 1.55×1.$5$, respectively, for every doubling of threads. At $64$ threads, NUMA affects the performance of our algorithms (NPS 4). In addition, increasing the number of threads makes the behavior of the algorithm less asynchronous, i.e., similar to the Jacobi iterative method, further contributing to the performance drop.
@@ 6. Conclusion
In this paper, we show how to adapt three approaches for community detection in dynamic graphs — Naive-dynamic (ND), Delta-screening (DS), and Dynamic Frontier (DF) — to a fast multicore implementation (Sahu et al., 2024a) of the Leiden algorithm (Traag et al., 2019). This enables users to track community changes over time while reducing the computational cost of updating communities as the graph evolves. Our experiments on a server equipped with a 64-core AMD EPYC-7742 CPU show that ND, DS, and DF Leiden achieve moderate speedups of 1.37×1.37$\times$1.37 ×, 1.47×1.47$\times$1.47 ×, and 1.98×1.98$\times$1.98 ×, respectively, on large graphs with randomized batch updates, compared to Static Leiden. These small gains are primarily because we need to run the algorithm for several passes, similar to Static Leiden, in order to obtain high quality top-level communities, and the reduction in runtime occurs only during the local-moving and refinement phases of the first pass of the algorithm. For every doubling of threads, our algorithms scale by 1.6×1.6$\times$1.6 ×. We hope these results encourage further investigation of dynamic Leiden algorithm.